# Carleman Estimates and Controllability of Forward Stochastic Parabolic Equations with General Dynamic Boundary Conditions

### v1

@@ 3 Observability Problems
$\begin{aligned} \textbf{I}_{4}&\leq C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}\psi||z_{\Gamma}||\nabla_{\Gamma}z_{\Gamma}|d\sigma dt+C\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z||\nabla_{\Gamma}z_{\Gamma}|d\sigma dt\\ &\quad\;+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial^{\mathcal{A}}_{\nu}\psi||z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial^{\mathcal{A}}_{\nu}\psi||z_{\Gamma}||F_{2}|d\sigma dt\\ &\quad\;+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial^{\mathcal{A}}_{\nu}\psi|z_{\Gamma}^{2}|\ell_{t}|d\sigma dt+2\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad\;+2\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z||F_{2}|d\sigma dt+2\mathbb{E}\iint_{\Sigma}\theta^{2}|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z||\ell_{t}|d\sigma dt.\end{aligned}$ (A.23)
Notice that $|\ell_{t}|\leq C\lambda Te^{2\mu|\psi|_{\infty}}\varphi^{2}$, and using Young’s inequality in the right-hand side of (A.23), we find that
$\begin{aligned} \textbf{I}_{4}&\leq C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt+C\mathbb{E}\iint_{\Sigma}\theta^{2}|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt+C\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad\;+C\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt+C\lambda^{2}\mu Te^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &\quad\;+CTe^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}z_{\Gamma}|^{2}d\sigma dt.\end{aligned}$ (A.24)
@@
$\textbf{I}_{5}=\begin{aligned} \mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}\widehat{Z}^{2}d\sigma dt=\mathbb{E}\iint_{\Sigma}\theta^{2}\left[-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}})\right]\widehat{Z}^{2}d\sigma dt,\end{aligned}$
@@
$\begin{aligned} \textbf{I}_{5}\leq C\lambda\mu T^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt+C\lambda Te^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (A.25)
Taking a large $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$, the inequality (A.25) implies hat
$\begin{aligned} \textbf{I}_{5}\leq C\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (A.26)
@@
$\textbf{I}_{6}\leq C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi z_{\Gamma}^{2}dxdt.$ (A.27)
@@
$\textbf{I}_{7}=-\mathbb{E}\iint_{\Sigma}\theta^{2}(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}d\sigma dt\leq C\mathbb{E}\iint_{\Sigma}\theta^{2}|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.$ (A.28)
@@
$\textbf{I}_{8}=2\mathbb{E}\iint_{\Sigma}\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})d\sigma dt\leq\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt+\mathbb{E}\iint_{\Sigma}\theta^{2}|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.$ (A.29)
@@
$\textbf{I}_{9}=2\mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt$ $=2\mathbb{E}\iint_{\Sigma}\theta^{2}\left[-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}})\right]z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt.$
@@
$\textbf{I}_{9}\leq C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt+C\lambda Te^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt,$
which implies that for large $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$,
$\textbf{I}_{9}\leq C\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt.$ (A.30)
Applying Young’s inequality for (A.30), we derive that
$\textbf{I}_{9}\leq C\lambda^{3}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+C\lambda\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt.$ (A.31)
Combining (A.13), (A.15), (A.27), (A.28), (A.19), (A.21), (A.22), (A.24), (A.26), (A.14), (A.29), (A.31), and taking a large enough $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$, we conclude that
$\begin{aligned} &2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dxdt+\mathbb{E}\iint_{Q}Bh^{2}dxdt\\ &+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt\\ &\quad\;+C\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt+C\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (A.32)
@@
From Lemma A.1, we have that for large $\mu\geq C$ and $\lambda\geq CT^{2}$,
$C\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt=C\mathbb{E}\iint_{Q}\theta^{2}\left[\lambda^{2}\mu^{2}\varphi^{2}\sum_{j,k=1}^{N}a^{jk}\psi_{x_{j}}\psi_{x_{k}}+\lambda\varphi O(\mu^{2})+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}})\right]Z^{2}dxdt.$
Taking a large $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$, it follows that
$C\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt\leq C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt.$ (A.33)
Since $\min_{x\in\overline{G\setminus G_{1}}}|\nabla\psi(x)|^{2}>0$, by Lemma A.1, we obtain for a large $\mu\geq C$ and $\lambda\geq CT^{2}$,
 $\begin{aligned} &\,2\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}\,dxdt+\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}Bh^{2}\,dxdt\\ &\geq 2\beta_{0}^{2}\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\Bigg[\left(\lambda\mu^{2}\varphi\min_{x\in\overline{G\setminus G_{1}}}|\nabla\psi|^{2}+\lambda\varphi O(\mu)\right)|\nabla h|^{2}\\ &\quad+\left(\lambda^{3}\mu^{4}\varphi^{3}\min_{x\in\overline{G\setminus G_{1}}}|\nabla\psi|^{4}+\lambda^{3}\varphi^{3}O(\mu^{4})+\lambda^{2}T\varphi^{3}O(\mu^{2}e^{2\mu|\psi|_{\infty}})\right)h^{2}\Bigg]\,dxdt.\end{aligned}$ (A.34)
From (A.34), we conclude for a large $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$,
 $\begin{aligned} &\,2\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}\,dxdt+\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}Bh^{2}\,dxdt\\ &\geq C\lambda\mu^{2}\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\varphi\left(|\nabla h|^{2}+\lambda^{2}\mu^{2}\varphi^{2}h^{2}\right)\,dxdt.\end{aligned}$ (A.35)
@@
$\frac{1}{C}\theta^{2}\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)\leq|\nabla h|^{2}+\lambda^{2}\mu^{2}\varphi^{2}h^{2}\leq C\theta^{2}\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right).$ (A.36)
@@
Combining (A.32), (A.33), (LABEL:5.232404) and (A.36), we conclude for large enough $\mu\geq C$ and $\lambda\geq C\left(e^{2\mu|\psi|_{\infty}}T+T^{2}\right)$,
@@
$\begin{aligned} &\,\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)dxdt\\ &+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &=\lambda\mu^{2}\mathbb{E}\int_{0}^{T}\left(\int_{G\setminus G_{1}}+\int_{G_{1}}\right)\theta^{2}\varphi\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)dxdt\\ &\quad\;+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt\\ &\quad\;+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt+\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt\\ &\quad\;+\lambda\mu^{2}\mathbb{E}\int_{0}^{T}\int_{G_{1}}\theta^{2}\varphi\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)dxdt.\end{aligned}$ (A.37)
@@
Finally, to eliminate the gradient term on the right-hand side of (A.37), we proceed exactly as in the proof of [27, Theorem 9.39] by choosing a cut-off function $\zeta\in C^{\infty}_{0}(G_{0};[0,1])$ with $\zeta\equiv 1$ in $G_{1}$, then computing $d(\theta^{2}\varphi\zeta^{2}z^{2})$ and integrating the equality over $Q$, and taking the expectation on both sides. We arrive at
@@
$\mathbb{E}\int_{0}^{T}\int_{G_{1}}\theta^{2}\varphi|\nabla z|^{2}\,dxdt\leq C\mathbb{E}\iint_{Q_{0}}\theta^{2}\left(\frac{1}{\lambda^{2}\mu^{2}}F_{1}^{2}+\lambda^{2}\mu^{2}\varphi^{3}z^{2}\right)dxdt.$ (A.38)

### v2

@@ 3 Intermediate Carleman Estimate without Divergence Terms
$\begin{aligned} \textbf{I}_{4}&\leq C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}\psi||z_{\Gamma}||\nabla_{\Gamma}z_{\Gamma}|d\sigma dt+C\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z||\nabla_{\Gamma}z_{\Gamma}|d\sigma dt\\ &\quad\;+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial^{\mathcal{A}}_{\nu}\psi||z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial^{\mathcal{A}}_{\nu}\psi||z_{\Gamma}||F_{2}|d\sigma dt\\ &\quad\;+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial^{\mathcal{A}}_{\nu}\psi|z_{\Gamma}^{2}|\ell_{t}|d\sigma dt+2\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad\;+2\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z||F_{2}|d\sigma dt+2\mathbb{E}\iint_{\Sigma}\theta^{2}|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z||\ell_{t}|d\sigma dt.\end{aligned}$ (3.25)
Notice that $|\ell_{t}|\leq C\lambda Te^{2\mu\|\psi\|_{\infty}}\varphi^{2}$, and using Young’s inequality in the right-hand side of (3.25), we find that
$\begin{aligned} \textbf{I}_{4}&\leq C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt+C\mathbb{E}\iint_{\Sigma}\theta^{2}|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt+C\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad\;+C\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt+C\lambda^{2}\mu Te^{2\mu\|\psi\|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &\quad\;+CTe^{2\mu\|\psi\|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}z_{\Gamma}|^{2}d\sigma dt.\end{aligned}$ (3.26)
@@
$\textbf{I}_{5}=\begin{aligned} \mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}\widehat{Z}^{2}d\sigma dt=\mathbb{E}\iint_{\Sigma}\theta^{2}\left[-\frac{2}{\beta_{0}}M\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu\|\psi\|_{\infty}})\right]\widehat{Z}^{2}d\sigma dt,\end{aligned}$
@@
$\begin{aligned} \textbf{I}_{5}\leq C\lambda\mu T^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt+C\lambda Te^{2\mu\|\psi\|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (3.27)
Taking a large $\lambda\geq C(e^{2\mu\|\psi\|_{\infty}}T+T^{2})$, the inequality (3.27) implies hat
$\begin{aligned} \textbf{I}_{5}\leq C\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (3.28)
@@
$\textbf{I}_{6}\leq C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi z_{\Gamma}^{2}dxdt.$ (3.29)
@@
$\textbf{I}_{7}=-\mathbb{E}\iint_{\Sigma}\theta^{2}(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}d\sigma dt\leq C\mathbb{E}\iint_{\Sigma}\theta^{2}|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.$ (3.30)
@@
$\textbf{I}_{8}=2\mathbb{E}\iint_{\Sigma}\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})d\sigma dt\leq\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt+\mathbb{E}\iint_{\Sigma}\theta^{2}|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.$ (3.31)
@@
$\textbf{I}_{9}=2\mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt$ $=2\mathbb{E}\iint_{\Sigma}\theta^{2}\left[-\frac{2}{\beta_{0}}M\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu\|\psi\|_{\infty}})\right]z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt.$
@@
$\textbf{I}_{9}\leq C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt+C\lambda Te^{2\mu\|\psi\|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt,$
which implies that for large $\lambda\geq C(e^{2\mu\|\psi\|_{\infty}}T+T^{2})$,
$\textbf{I}_{9}\leq C\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt.$ (3.32)
Applying Young’s inequality for (3.32), we derive that
$\textbf{I}_{9}\leq C\lambda^{3}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+C\lambda\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt.$ (3.33)
Combining (3.15), (3.17), (3.29), (3.30), (3.21), (3.23), (3.24), (3.26), (3.28), (3.16), (3.31), (3.33), and taking a large enough $\mu\geq C$ and $\lambda\geq C(e^{2\mu\|\psi\|_{\infty}}T+T^{2})$, we conclude that
$\begin{aligned} &2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dxdt+\mathbb{E}\iint_{Q}Bh^{2}dxdt\\ &+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt\\ &\quad\;+C\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt+C\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (3.34)
@@
From Lemma 3.1, we have that for large $\mu\geq C$ and $\lambda\geq CT^{2}$,
$C\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt=C\mathbb{E}\iint_{Q}\theta^{2}\left[\lambda^{2}\mu^{2}\varphi^{2}\sum_{j,k=1}^{N}a^{jk}\psi_{x_{j}}\psi_{x_{k}}+\lambda\varphi O(\mu^{2})+\lambda T\varphi^{2}O(e^{2\mu\|\psi\|_{\infty}})\right]Z^{2}dxdt.$
Taking a large $\lambda\geq C(e^{2\mu\|\psi\|_{\infty}}T+T^{2})$, it follows that
$C\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt\leq C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt.$ (3.35)
Since $\min_{x\in\overline{G\setminus G_{1}}}|\nabla\psi(x)|^{2}>0$, by Lemma 3.1, we obtain for a large $\mu\geq C$ and $\lambda\geq CT^{2}$,
 $\begin{aligned} &\,2\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}\,dxdt+\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}Bh^{2}\,dxdt\\ &\geq 2\beta_{0}^{2}\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\bigg[\left(\lambda\mu^{2}\varphi\min_{x\in\overline{G\setminus G_{1}}}|\nabla\psi|^{2}+\lambda\varphi O(\mu)\right)|\nabla h|^{2}\\ &\quad+\left(\lambda^{3}\mu^{4}\varphi^{3}\min_{x\in\overline{G\setminus G_{1}}}|\nabla\psi|^{4}+\lambda^{3}\varphi^{3}O(\mu^{4})+\lambda^{2}T\varphi^{3}O(\mu^{2}e^{2\mu\|\psi\|_{\infty}})\right)h^{2}\bigg]\,dxdt.\end{aligned}$ (3.36)
From (3.36), we conclude for a large $\mu\geq C$ and $\lambda\geq C(e^{2\mu\|\psi\|_{\infty}}T+T^{2})$,
 $\begin{aligned} &\,2\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}\,dxdt+\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}Bh^{2}\,dxdt\\ &\geq C\lambda\mu^{2}\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\varphi\left(|\nabla h|^{2}+\lambda^{2}\mu^{2}\varphi^{2}h^{2}\right)\,dxdt.\end{aligned}$ (3.37)
@@
$\frac{1}{C}\theta^{2}\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)\leq|\nabla h|^{2}+\lambda^{2}\mu^{2}\varphi^{2}h^{2}\leq C\theta^{2}\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right).$ (3.38)
@@
Combining (3.34), (3.35), (LABEL:5.232404) and (3.38), we conclude for large enough $\mu\geq C$ and $\lambda\geq C\left(e^{2\mu\|\psi\|_{\infty}}T+T^{2}\right)$,
@@
$\begin{aligned} &\,\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)dxdt\\ &+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &=\lambda\mu^{2}\mathbb{E}\int_{0}^{T}\left(\int_{G\setminus G_{1}}+\int_{G_{1}}\right)\theta^{2}\varphi\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)dxdt\\ &\quad\;+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt\\ &\quad\;+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt+\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt\\ &\quad\;+\lambda\mu^{2}\mathbb{E}\int_{0}^{T}\int_{G_{1}}\theta^{2}\varphi\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)dxdt.\end{aligned}$ (3.39)
@@
Finally, to eliminate the gradient term on the right-hand side of (3.39), we proceed exactly as in the proof of [34, Theorem 9.39] by choosing a cut-off function $\zeta\in C^{\infty}_{0}(\mathcal{B};[0,1])$ with $\zeta\equiv 1$ in $G_{1}$, then computing $d(\theta^{2}\varphi\zeta^{2}z^{2})$ and integrating the equality over $Q$, and taking the expectation on both sides. We arrive at
@@
$\mathbb{E}\int_{0}^{T}\int_{G_{1}}\theta^{2}\varphi|\nabla z|^{2}\,dxdt\leq C\mathbb{E}\int_{0}^{T}\int_{\mathcal{B}}\theta^{2}\left(\frac{1}{\lambda^{2}\mu^{2}}F_{1}^{2}+\lambda^{2}\mu^{2}\varphi^{3}z^{2}\right)dxdt.$ (3.40)